Suppose you deposit $2,655.00 into an account today. In 10.00 years the account is worth $3,994.00. The account earned ____% per year.

Answer Format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))
Answer #1

We use the formula

A=P(1+r/100)^n

Where A=future value

P=present value

R is rate of interest

N=time period

3994=2655*(1+r/100)^10

(3994/2655)^(1/10)=1+r/100

1+r/100=1.0417

Hence r=(1.0417-1)*100

=4.17%(approx).